# Listen to Look into the Future: Audio-Visual Egocentric Gaze Anticipation

### 07 Dec 2023 v2

@@ 3.3 Contrastive Learning for Audio-Visual Fusion
In addition to using KL-Divergence loss to supervise gaze anticipation, we propose to leverage the intrinsic alignment of visual and audio modalities to learn a more robust audio-visual representation by using a contrastive learning scheme. Multi-modal contrastive loss has been proved to be effective in self-supervised learning [3, 2]. Rather than calculating the contrastive loss directly on embedded features, we propose to use the reweighted video and audio representations from the spatial and temporal fusion modules. In our experiments, we show this is a more effective representation learning method for egocentric gaze anticipation.
To this end, we reweight the raw audio embedding $\psi(a)\in\mathbb{R}^{T\times M\times D}$ from the audio encoder by temporal weights $u_{a,t}$ from the temporal fusion module in a similar way to Eq. 7. We then get the reweighted audio feature as:
$\small u_{a}=\psi(a)\otimes u_{a,t}\in\mathbb{R}^{T\times M\times D}$ (8)
We don’t use an additional learnable token to aggregate information from other tokens as prior works did [3, 2, 36]. We instead average all tokens of ${u_{v}}$ and ${u_{a}}$ respectively to obtain the single-vector representations $u^{\prime}_{v},u^{\prime}_{a}\in\mathbb{R}^{1\times D}$ and then map them to a common space using linear layers followed by L2-norm. It can be formulated as $w_{v}=Norm\left(f_{1}(u^{\prime}_{v})\right)$ and $w_{a}=Norm\left(f_{2}(u^{\prime}_{a})\right)$, where $f_{1}(\cdot),f_{2}(\cdot)$ are linear layers. The resulting visual vector and audio vector are denoted as $w_{v},w_{a}\in\mathbb{R}^{1\times D^{\prime}}$, where $D^{\prime}$ is the new dimension of the common space. Within each mini-batch, corresponding audio and visual embeddings are considered as positive pairs, and all other pairwise combinations are considered as negative pairs. Following [36], we calculate video-to-audio loss and audio-to-video loss separately. The video-to-audio contrastive loss is defined as
$\small\mathcal{L}^{v2a}_{cntr}=-\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\log\frac{\exp({w_{v}^{(i)}}^{T}w_{a}^{(i)}/\mathcal{T})}{\sum_{j\in\mathcal{B}}\exp({w_{v}^{(i)}}^{T}w_{a}^{(j)}/\mathcal{T})},$ (9)
where $\mathcal{B}$ is the training batch =$\mathcal{B}=\{1,2,\dots,n\}$ and $\mathcal{T}$ is the temperature factor. Superscripts $(i)$ and $(j)$ denote the $i$-th and $j$-th samples in the batch. The audio-to-video loss is defined in a symmetric way. Finally, the contrastive loss is defined as $\mathcal{L}_{cntr}=\mathcal{L}_{cntr}^{v2a}+\mathcal{L}_{cntr}^{a2v}$. $\mathcal{L}_{kld}$ and $\mathcal{L}_{cntr}$ are linearly combined with a parameter $\alpha$ for the final training loss, i.e. $\mathcal{L}=\mathcal{L}_{kld}+\alpha\mathcal{L}_{cntr}$.
@@ 3.4 Implementation Details
In our experiments, we set observation time $\tau_{o}$ as $3$ seconds and anticipation time $\tau_{a}$ as $2$ seconds. For video input, we sample 8 frames from the observable segment and resize to a spatial size of 256$\times$256. For audio input, following [31], we first resample the audio signal to 24kHz and set a time window with $\Delta t_{w}=1.28s$ to crop the audio segment corresponding to each video frame. We then convert it to a log-spectrogram using a STFT with window size 10ms and hop length 5ms. The number of frequency bands is set as 256 resulting in a spectrogram matrix of size 256$\times$256. The output of the decoder is the gaze distribution on 8 frames uniformly sampled from the 2-second anticipation time. More details about model architecture and training hyper-parameters can be found in the supplementary.
@@ 4 Experiments
We first introduce the datasets and evaluation metrics used in our experiments. We then present detailed ablation studies to validate the contribution of each component in our method, and demonstrate the performance improvement over prior SOTA methods for gaze anticipation as well as gaze estimation models applied to the gaze anticipation task. Finally, we visualize the predictions and weights of our model to provide qualitative insight into our method.
@@ 4.1 Experiment Setup
Datasets. We conduct experiments on two egocentric datasets [^1] Note that another widely used gaze estimation benchmark EGTEA Gaze+ [35] did not release audio data and thus is not feasible for our study. that contain aligned video and audio streams and gaze tracking data – Ego4D [17] and Aria [29]. The Ego4D eye-tracking subset is collected in social settings and totals 31 hours of egocentric videos from 80 participants. All videos have a fixed 30 fps frame rate and spatial resolution of 1088$\times$1080, and audio streams are recorded with a sampling rate of 44.1kHz. We use the train/test split released in [33] in our experiments, i.e., 15310 video segments for training and the other 5202 video segments for testing.
@@
The Aria dataset contains 143 egocentric videos (totaling 7.5 hours) collected with Project Aria glasses. It covers a variety of everyday activities including cooking, exercising and spending time with friends. All videos have a fixed 30 fps frame rate and spatial resolution of 1408$\times$1408. A sliding window is used to trim long videos into 5-second video segment with a stride of 2 seconds. We use 107 videos (10456 segments) for training and 36 videos (2901 segments) for testing. We will release our split to facilitate future studies in this direction.
@@
Evaluation Metrics. As suggested in recent work on egocentric gaze estimation [33], AUC score can easily get saturated due to the long-tailed distribution of gaze on 2D video frames. Therefore, we follow [33, 35] to adopt F1 score, recall and precision as our evaluation metrics.
4.2 Experimental Results
4.2.1 Ablation Study
@@
We first quantify the performance contribution of each key module from our proposed method. Specifically, we denote the model with only spatial fusion module as S-fusion, the model with only temporal fusion as T-fusion, the model uses our proposed spatial-temporal fusion module without the contrastive learning schema as STS. As demonstrated in Table 1, compared with models trained solely on RGB frames (Vision only), S-fusion and T-fusion boost the F1 score by +1.4% and +1.5% on Ego4D, and +1.1% and +1.1% on Aria. Moreover, the STS model further achieves a F1 score of 39.2% on Ego4D and 59.3% on Aria. These results suggest that both the spatial and and the temporal correlation between video and audio signal play a vital role for egocentric gaze anticipation. Contrastive loss further improves F1 by +0.5% and +0.6% suggesting its notable contributions to audio-visual representative learning.
@@
F1 Rec. Prec. F1 Rec. Prec.
@@
Table 1: Ablations on each key component of our proposed model. The best results are highlighted with boldface. See Section 4.2.1 for further discussion.
4.2.2 Analysis on Fusion and Contrastive Learning
@@
F1 Rec. Prec. F1 Rec. Prec.
@@
Table 2: Analysis on proposed fusion strategies. The best results are highlighted with boldface. STS refers to the proposed spatial-temporal separable fusion method (without contrastive learning). See Section 4.2.2 for details.
Directly feeding all visual and audio tokens into a fusion layer (i.e. joint fusion) is a widely used approach for audio-visual saliency prediction [58, 62, 6] and sound source localization [21, 47, 5, 7]. To show the superiority of the proposed spatial-temporal separable (STS) fusion approach, we provide additional analysis of three joint fusion strategies: (1) concatenating audio and visual embeddings along channel dimension (denoted as Concat.) as in [31, 27]; (2) feeding all embedded video and audio tokens together into a standard self-attention layer (denoted as Vanilla SA), which is inspired by [37, 62]; and (3) a variant of STS that connects the two fusion modules sequentially, i.e. feeding embeddings to the temporal fusion module first and then feeding its output to the spatial fusion module (denoted as Seq. ST). We further detail how our design differs from the above fusion strategies in the supplementary.
As shown in Table 2, Concat. and Vanilla SA has limited improvement over the vision-only baseline, suggesting that previous audio-visual fusion methods, which follow these strategies, are sub-optimal for our problem setting. In contrast, our proposed fusion strategy yields larger performance boost (+2.0% on Ego4D and +1.8% on Aria), even without using the contrastive loss (STS), which shows the benfits of our proposed spatial-temporal separable fusion mechanism. Additionally, our model also consistently surpasses its sequential variant (Seq. ST) by +0.7% on both datasets, further demonstrating the benefits of explicitly reasoning spatial and temporal correlation separately.
@@
F1 Rec. Prec. F1 Rec. Prec.
@@
Table 3: Analysis on proposed contrastive learning schema. The best results are highlighted with boldface. The highlighted row refers to the complete CSTS model. See Section 4.2.2 for details.
We also evaluate the benefits of our contrastive learning scheme in Table 3. Here, we consider another baseline (denoted as Vanilla Contr) that calculates the contrastive loss using raw video and audio embeddings (i.e. $\phi(x)$ and $\psi(a)$), as is typical in prior work [56, 39, 16, 18]. Our strategy of adding contrastive loss on fused features is denoted as Post Contr. Vanilla Contr makes only minor differences on Vanilla SA model and even slightly reduces performance when accompanied by our proposed STS mechanism. In contrast, our proposed Post Contr schema improves the performance of Vanilla SA by +0.4% and 0.8% and improves STS by +0.5% and +0.6% on the two datasets. These results further suggest that post-fusion contrastive learning is more robust for audio-visual representation learning. Additional experiments with different contrastive learning strategies are provided in the supplementary.
4.2.3 Comparison with State-of-the-art Methods
Figure 3: The performance of gaze anticipation in each frame. Our model consistently outperforms all prior methods by a notable margin.
Most existing works on egocentric gaze modeling target egocentric gaze estimation rather than anticipation. In order to provide a thorough comparison, in addition to comparing against SOTA egocentric gaze anticipation models (DFG[65], DFG+[66]), we also adapt the recent SOTA egocentric gaze estimation model GLC[33] and all baselines from [33] (I3D-Res50 [61], MViT[12], GazeMLE[35] and AttnTrans[24]) to the anticipation task. Detailed results are summarized in Table 4.
@@ 4.2.3 Comparison with State-of-the-art Methods
F1 Rec. Prec. F1 Rec. Prec.
@@
Table 4: Comparison with previous state-of-the-art models on egocentric gaze anticipation. We also adapt previous egocentric gaze estimation approaches to the anticipation setting for a more thorough comparison. The best results are highlighted with boldface. See Section 4.2.3 for further discussion.
Our method outperforms its direct competitor DFG+, which is the previous SOTA model for egocentric gaze anticipation, by +2.4% F1 on Ego4D and +2.3% F1 on Aria. Note that the original DFG and DFG+ used a less powerful backbone encoder, so for fair comparison, we reimplement their method using the same MViT backbone as our method. We also observe that methods originally designed for egocentric gaze estimation still work as strong baselines for the egocentric gaze anticipation task. Our proposed CSTS model also outperforms these methods, surpassing the recent SOTA for egocentric gaze estimation, GLC, by +1.9% F1 on Ego4D and +1.6% F1 on Aria.
In addition, we evaluate gaze anticipation on each anticipation time step independently and compare with previous methods in Fig. 3. Unsurprisingly, the anticipation problem becomes more challenging as the anticipation time step increases farther into the future. Our CSTS method consistently outperforms all baselines at all future time steps. We also note that our model produces new SOTA results on egocentric gaze estimation, demonstrating the generalizability and robustness of our approach across gaze modeling tasks. We include these results in the supplementary.
@@
4.2.4 Visualization of Predictions
We visually showcase the anticipation results of CSTS and the baselines in Fig. 5. We can see that GazeMLE [35] and AttnTransit [24] produce more ambiguity in prediction heatmaps. Other methods fail to anticipate the true gaze target, and are likely misled by other salient objects. Our CSTS approach produces the best gaze anticipation results among all methods. We attribute this improvement to our novel model design that effectively leverages both audio and visual cues for forecasting the gaze targets.
4.2.5 Visualization of Learned Correlations
We provide further insight on our model by visualizing the audio-visual correlation from the spatial fusion module. For each time step $t$, we calculate the correlation of each visual token with the single audio token and map it back to the input frames. The correlation heatmaps are shown in Fig. 4. In the first example, the speaker in the middle speaks, then turns her head around to talk with a social partner in the background (e.g. frame 1-3). We observe that our model captures that the audio signal has the highest correlation with spatial region of the speaker while she is speaking. Then, when she stops talking and turns her head back, the correlation is highest in the background regions, indicating the potential location of her social partner. The second example illustrates a similar phenomenon; the model captures the speaker at the beginning when she is talking, then attend to background locations when she stops. These examples suggest our model has the capability to model the audio-visual correlations in spatial dimension to learn a robust audio-visual representation.

### 22 Mar 2024 v3

@@ 3.3 Contrastive Learning for Audio-Visual Fusion
In addition to using KL-Divergence loss to supervise gaze anticipation, we propose to leverage the intrinsic alignment of visual and audio modalities to learn a more robust audio-visual representation by using a contrastive learning scheme. Multi-modal contrastive loss has been proved to be effective in self-supervised learning [3, 2]. Rather than calculating the contrastive loss directly on the raw embedded features, we innovatively propose to use the reweighted video and audio representations from the spatial and temporal fusion modules, which has not been studied in prior works. In our experiments, we show this is a more effective representation learning method for egocentric gaze anticipation.
To this end, we reweight the raw audio embedding $\psi(a)\in\mathbb{R}^{T\times M\times D}$ from the audio encoder by temporal weights $u_{a,t}$ from the temporal fusion module in a similar way to Eq. 7. We then get the reweighted audio feature as
$u_{a}=\psi(a)\otimes u_{a,t}\in\mathbb{R}^{T\times M\times D}$ (8)
We don’t use an additional learnable token to aggregate information from other tokens as prior works did [3, 2, 39]. We instead average all tokens of ${u_{v}}$ and ${u_{a}}$ respectively to obtain the single-vector representations $u^{\prime}_{v},u^{\prime}_{a}\in\mathbb{R}^{1\times D}$ and then map them to a low-dimensional common space using linear layers followed by L2 normalization. It can be formulated as $w_{v}=Norm\left(f_{1}(u^{\prime}_{v})\right)$ and $w_{a}=Norm\left(f_{2}(u^{\prime}_{a})\right)$, where $f_{1}(\cdot),f_{2}(\cdot)$ are linear layers. The resulting visual vector and audio vector are denoted as $w_{v},w_{a}\in\mathbb{R}^{1\times D^{\prime}}$, where $D^{\prime}$ is the new dimension of the common space. Within each mini-batch, corresponding audio and visual embeddings are considered as positive pairs, and all other pairwise combinations are considered as negative. Following [39], we calculate video-to-audio loss and audio-to-video loss separately. The video-to-audio contrastive loss is defined as
$\mathcal{L}^{v2a}_{cntr}=-\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\log\frac{\exp({w_{v}^{(i)}}^{T}w_{a}^{(i)}/\mathcal{T})}{\sum_{j\in\mathcal{B}}\exp({w_{v}^{(i)}}^{T}w_{a}^{(j)}/\mathcal{T})},$ (9)
where $\mathcal{B}$ is the training batch =$\mathcal{B}=\{1,2,\dots,n\}$ and $\mathcal{T}$ is the temperature factor. Superscripts $(i)$ and $(j)$ denote the $i$-th and $j$-th samples in the batch. The audio-to-video loss is defined in a symmetric way. Finally, the contrastive loss is defined as $\mathcal{L}_{cntr}=\mathcal{L}_{cntr}^{v2a}+\mathcal{L}_{cntr}^{a2v}$. $\mathcal{L}_{kld}$ and $\mathcal{L}_{cntr}$ are linearly combined with a parameter $\alpha$ for the final training loss, i.e., $\mathcal{L}=\mathcal{L}_{kld}+\alpha\mathcal{L}_{cntr}$.
@@ 3.4 Implementation Details
In our experiments, we set observation time $\tau_{o}$ as $3$ seconds and anticipation time $\tau_{a}$ as $2$ seconds. For video input, we sample 8 frames from the observable segment and resize to a spatial size of 256$\times$256. For audio input, following [34], we first resample the audio signal to 24kHz and set a time window with $\Delta t_{w}=1.28s$ to crop the audio segment corresponding to each video frame. We then convert it to a log-spectrogram using a STFT with window size 10ms and hop length 5ms. The number of frequency bands is set as 256 resulting in a spectrogram matrix of size 256$\times$256. The output of the decoder is the gaze distribution on 8 frames uniformly sampled from the 2-second anticipation time. More details about model architecture and training hyper-parameters can be found in supplementary.
@@ 4 Experiments
We first introduce the datasets and evaluation metrics used in our experiments. We then present detailed ablation studies to validate the contribution of each component in our method, and demonstrate the performance improvement over prior state-of-the-art methods for gaze anticipation as well as gaze estimation models applied to the gaze anticipation task. Finally, we visualize the predictions and weights of our model to provide qualitative insight into our method.
@@ 4.1 Experiment Setup
Datasets. We conduct experiments on two egocentric datasets that contain aligned video and audio streams and gaze tracking data – Ego4D [20] and Aria [43]. Note that another widely used gaze estimation benchmark EGTEA Gaze+ [38] does not release audio data and thus is not feasible for our study. Other popular egocentric video datasets, such as Epic-Kitchens [11] and Charades-Ego [62], are also not applicable to our task because they don’t have eye-tracking labels. Ego4D and Aria are the two largest public datasets that provide all necessary data and labels (i.e., egocentric videos, aligned audio streams and eye-tracking data) for egocentric audio-visual gaze anticipation.
The Ego4D eye-tracking subset is collected in social settings (i.e., social interaction benchmark) and totals 31 hours of egocentric videos from 80 participants. All videos have a fixed 30 fps frame rate and spatial resolution of 1088$\times$1080, and audio streams are recorded with a sampling rate of 44.1kHz. We use the train/test split released in [36] in our experiments, i.e., 15310 video segments for training and the other 5202 video segments for testing.
@@
The Aria dataset contains 143 egocentric videos (totaling 7.3 hours) collected with Project Aria glasses. It covers a variety of indoor everyday activities including cooking, exercising and spending time with friends. All videos have a fixed 20 fps frame rate and spatial resolution of 1408$\times$1408. A sliding window is used to trim long videos into 5-second video segment with a stride of 2 seconds. We use 107 videos (10456 segments) for training and 36 videos (2901 segments) for testing. We will release our split to facilitate future studies in this direction.
@@
Evaluation Metrics. As suggested in recent work on egocentric gaze estimation [36], AUC score can easily get saturated due to the long-tailed distribution of gaze on 2D video frames. Therefore, we follow [36, 38] to adopt F1 score (primary), recall and precision as our evaluation metrics.
4.2 Ablation Study
@@
We first quantify the performance contribution of each key module from our proposed method. Specifically, we denote the model only using our proposed spatial fusion module as S-fusion, the model only using our proposed temporal fusion module as T-fusion, the model using both modules and our spatial-temporal separable fusion strategy without the contrastive learning schema as STS. We finally present the performance of our full CSTS model (i.e., STS + contrastive learning). As demonstrated in Tab. 1, compared with models trained solely on RGB frames (Vision only), S-fusion and T-fusion boost the F1 score by +1.4% and +1.5% on Ego4D, and +1.1% and +1.1% on Aria. Moreover, the STS model further achieves a F1 score of 39.2% on Ego4D and 59.3% on Aria. These results suggest that both the spatial and and the temporal correlation between video and audio signal play a vital role for egocentric gaze anticipation. Contrastive loss further improves F1 score by +0.5% and +0.6% suggesting its contributions to audio-visual representative learning. We also observe that the full model doesn’t achieve the best in recall. This is because some incomplete baselines don’t leverage audio modality as effectively as the full model and thus produce more uncertainty in output, resulting in higher recall and lower precision. Therefore, we consider F1 as the primary metric. Similar phenomenon is also observed in the following experiments.
Table 1: Ablations on each key component of our proposed model. CSTS (highlighted in green) refers to the complete model of our approach. The best results are highlighted with boldface. Please refer to Sec. 4.2 for more discussions.
@@
F1 Score Recall Precision F1 Score Recall Precision
@@
Table 2: Analysis on proposed fusion strategies. The best results are highlighted with boldface. STS (highlighted in green) refers to the proposed spatial-temporal separable fusion method (without contrastive learning). More discussions are in Sec. 4.3.
@@
F1 Score Recall Precision F1 Score Recall Precision
@@
4.3 Analysis on Fusion and Contrastive Learning Strategies
Directly feeding all visual and audio tokens into a fusion layer (i.e., joint fusion) is a widely used approach for audio-visual saliency prediction [67, 73, 6] and action recognition [34, 17, 70]. To show the superiority of the proposed spatial-temporal separable (STS) fusion approach in handling the unique challenges of our task, we provide additional comparison with four joint fusion strategies that are widely used in audio-visual saliency prediction and audio-visual action recognition. Specifically, the four strategies are (1) fusing two modalities with a few linear layers [17] (denoted as Linear); (2) feeding video and audio embeddings to a single bilinear layer [31, 75] (denoted as Bilinear); (3) concatenating audio and visual embeddings along channel dimension (denoted as Concat.) as in [34, 31]; (4) feeding all embedded video and audio tokens together into a standard self-attention layer (denoted as Vanilla SA), inspired by [40, 73]. We replace our fusion modules with the four strategies in our framework for a fair comparison. We elaborate the implementation details of each baseline in supplementary.
As shown in Tab. 2, Linear, Bilinear, Concat. and Vanilla SA methods have limited improvement over the vision-only baseline, suggesting that previous fusion strategies for audio-visual saliency prediction and general action recognition are sub-optimal for our problem setting. In contrast, our proposed fusion strategy (STS) yields larger performance boost (+2.0% on Ego4D and +1.8% on Aria) even without using the contrastive loss, which shows the benefits of spatial-temporal separable fusion mechanism. The possible reason is that prior joint fusion methods are designed for third-person videos without a drastic viewpoint change. However, forecasting gaze in egocentric view has the unique challenges caused by camera movement and the latency of gaze response to audio stimuli. Our approach fuses two modalities in space and time separately and hence avoids spurious correlations that may happen in joint fusion baselines.
Table 3: Analysis on the proposed contrastive learning schema. Post Contr denotes our post-fusion contrastive learning. STS + Post Contr refers to the complete CSTS model. The best results are highlighted with boldface. More discussions are in Sec. 4.3.
@@ 4.3 Analysis on Fusion and Contrastive Learning Strategies
F1 Score Recall Precision F1 Score Recall Precision
@@
We also evaluate the benefits of our proposed post-fusion contrastive learning scheme in Tab. 3. Here, we consider another baseline (denoted as Vanilla Contr) that calculates the contrastive loss using raw video and audio embeddings (i.e., $\phi(x)$ and $\psi(a)$ in Fig. 2), as is typical in prior work [65, 44, 19, 21]. Our novel strategy of adding contrastive loss on fused features is denoted as Post Contr. Vanilla Contr makes only minor differences on Vanilla SA model and even slightly reduces performance when accompanied by our proposed STS mechanism. In contrast, our proposed Post Contr scheme improves the performance of Vanilla SA by +0.4% and 0.8% and improves STS by +0.5% and +0.6% on the two datasets. These results further suggest that post-fusion contrastive learning is more robust for audio-visual learning in our task. More experiments of different contrastive learning strategies are provided in supplementary.
4.4 Comparison with State-of-the-art Methods
Table 4: Comparison with state-of-the-art models on egocentric gaze anticipation. We also adapt previous egocentric gaze estimation approaches to the anticipation setting for a more thorough comparison. The best results are highlighted with boldface. The green row shows our model performance. Please refer to Sec. 4.4 for more discussions.
@@ 4.4 Comparison with State-of-the-art Methods
F1 Score Recall Precision F1 Score Recall Precision
@@
Figure 3: The performance of gaze anticipation in each frame. Our model (CSTS) consistently outperforms all prior methods by a notable margin.
Most existing works on egocentric gaze modeling target at egocentric gaze estimation rather than anticipation. In order to provide a thorough comparison, in addition to comparing against SOTA egocentric gaze anticipation models (DFG[77], DFG+[76]), we also adapt the recent SOTA egocentric gaze estimation model GLC[36] and all baselines from [36] (I3D-Res50 [71], MViT[14], GazeMLE[38] and AttnTrans[29]) to the anticipation task.
As presented in Tab. 4, our method outperforms its direct competitor DFG+, which is the previous SOTA model for egocentric gaze anticipation, by +2.4% F1 on Ego4D and +2.3% F1 on Aria. Note that the original DFG and DFG+ used a less powerful backbone encoder, so for fair comparison, we reimplement their method using the same MViT backbone as our method. We also observe that methods originally designed for egocentric gaze estimation still work as strong baselines for the egocentric gaze anticipation task. Our proposed CSTS model also outperforms these methods, surpassing the recent SOTA for egocentric gaze estimation – GLC by +1.9% F1 on Ego4D and +1.6% F1 on Aria. In addition, We also incorporate audio stream into the strongest baseline (GLC) by a straightforward concatenation whose F1 score is 38.1% on Eg4D and 58.5% on Aria. The marginal gain over GLC (+0.3%/+0.2%) suggests that simply using audio stream in a strong baseline without specific design leads to sub-optimal solution in egocentric gaze anticipation problem, which in turn validates the effectiveness and necessity of our approach.
In addition, we evaluate gaze anticipation on each anticipation time step independently and compare with previous methods in Fig. 3. Unsurprisingly, the anticipation problem becomes more challenging as the anticipation time step increases farther into the future. Our CSTS method consistently outperforms all baselines at all future time steps. Moreover, we note that our model also produces new SOTA results on egocentric gaze estimation, demonstrating the generalizability and robustness of our approach across gaze modeling tasks. We include these results in supplementary.
@@
4.5 Visualization of Predictions
We visually showcase the anticipation results of CSTS and the baselines in Fig. 4. We can see that GazeMLE [38] and AttnTransit [29] produce more uncertainty in prediction heatmaps. Other methods fail to anticipate the true gaze target, and are likely to be misled by other salient objects. Our CSTS approach produces the best gaze anticipation results among all methods. We attribute this improvement to our novel model design that effectively addresses the unique challenges of forecasting gaze targets in egocentric view.
4.6 Visualization of Learned Correlations
We provide further insight on our model by visualizing the audio-visual correlations from the spatial fusion module. For each time step $t$, we calculate the correlation of each visual token with the single audio token and map it back to the input frames. The correlation heatmaps are shown in Fig. 5. In the first example, the speaker in the middle speaks, then turns her head around to talk with a social partner in the background (frame 1-3). We observe that our model captures that the audio signal has the highest correlation with spatial region of the speaker while she is speaking. Then, when she stops talking and turns her head back, the correlation is highest in the background regions, indicating the potential location of her social partner. The second example illustrates a similar phenomenon: the model captures the speaker at the beginning when she is talking, then attends to background locations when she stops. These examples suggest our model has the capability to model the audio-visual correlations in spatial dimension to learn a robust audio-visual representation.